OpenStudy (anonymous):

The length of a rectangle is 3 inches more than its width. The area of the rectangle is 40 square inches. What is the length, in inches, of the rectangle? (1) 5 (3) 8.5 (2) 8 (4) 11.5

OpenStudy (anonymous):

Set L = W + 3 Area equals L * W So (W+3)*W = 40 (put everything on one side) w^2 +3w -40 = 0 (factor) (w+8)(w-5) = 0 W = 5, -8 Since W can't be negative, W = 5 If W =5, Then L = W + 3 = 8 and A = 5 * 8 = 40